This is statistics but im hoping you can help...
The probability of hilary running for president in 2008 is 45%, the probability for george running for president is 20%, they are both independent, what is the probability hilary runs and george doesnt?
a. 9% b.18% c.25% d.36% e.45%


a survey recorded gender and wether the student was left or right handed. if handedness and gender are independent, how many were lefty girls? total boys=66 total girls=54 total lefts=20 total rights=100
a.4 b.7 c.9 d.10


a council has 6 men and 3 women, randomly select 2 of them to co-chair a new committee whats the probability the people are the same gender?
a. 4/9 b. 1/2 c. 5/9 d. 5/8 e. 7/8


12% of ice cream cones sold are jumbo size. whats the probabilty that the first jumbo cone sold is the fourth one sold?
a. 8% b.33% c.40% d.60% e. none of these are right

looking at the above, whats the probability that there are 2 jumbos sold out of the frst 6 cones sold?
a. 11% b. 23% c. 13% d.97% e.none of these are right


you toss a coin 20 times, 15 times it lands on heads, if you flip 180 more times, how many heads will end up from the 200 total flips?
a. 100 b.105 c. 110 d.115



whats true of 2 random variables if we are to add to their variances?
a.must be disjoint
b.must be independent
c. must be both
d. must be neither


police say 40% accidents involve speeding, 25% beer, and 10% both speeding and beer
whats the probability an accident involved neither?
are they independent?


i know its a lot but i was sick for a lot of days and now my teacher isnt here so im just trying to find problems to catch up on, i dont mean to throw it all on you but i just dont understand this at all, if you know any of these answers it would help so much!
thanks!

Answer by stanbon(48568)

The probability of hilary running for president in 2008 is 45%, the probability for george running for president is 20%, they are both independent, what is the probability hilary runs and george doesnt?
a. 9% b.18% c.25% d.36% e.45%
If events A and B are independent P(A and B)=P(A)*P(B)
P(H and not G) = P(H)*P(not G) = 0.45 * 0.80 = 0.36
--------------------

a survey recorded gender and wether the student was left or right handed. if handedness and gender are independent, how many were lefty girls? total boys=66 total girls=54 total lefts=20 total rights=100
a.4 b.7 c.9 d.10
Draw a contingency table with rows left and right and columns boys and girls
Put the numbers in the correct boxes. Try each of the possible answers in
the left-girls box; figure out what the other box numbers would have to be;
if you get a contradiction when you try a particular possible answer, that
answer is wrong. I find that both 4 and 10 are possible answers.
================

a council has 6 men and 3 women, randomly select 2 of them to co-chair a new committee whats the probability the people are the same gender?
a. 4/9 b. 1/2 c. 5/9 d. 5/8 e. 7/8
Number of pairs that are same gender = 6C2 + 3C2 = 15+3 = 18
Number of unrestricted pairs = 9C2 = 36
Prob(people are same gender) = 18/36 = 1/2
-------------------

12% of ice cream cones sold are jumbo size. whats the probabilty that the first jumbo cone sold is the fourth one sold?
a. 8% b.33% c.40% d.60% e. none of these are right
You need to pick (not jumbo) three times, then pick a jumbo.
P(NNNJ)= (0.88)^3*(0.12) =0.0818
===================
looking at the above, whats the probability that there are 2 jumbos sold out of the frst 6 cones sold?
a. 11% b. 23% c. 13% d.97% e.none of these are right
This is a binomial problem with n=6,p=0.12,x=2
P(X=2) = 6C2(0.12)^2(0.88)^4 = 0.1295...
------------

you toss a coin 20 times, 15 times it lands on heads, if you flip 180 more times, how many heads will end up from the 200 total flips?
a. 100 b.105 c. 110 d.115
Based on the 20 throw results you have a biased coin where P(H)=3/4
So (3/4)*180 = 135
Total number of heads out of 200 will be 150.
==============
whats true of 2 random variables if we are to add to their variances?
a.must be disjoint
b.must be independent
c. must be both
d. must be neither
I don't understand your question.
====================

police say 40% accidents involve speeding, 25% beer, and 10% both speeding and beer
whats the probability an accident involved neither?
are they independent?
P(speed or beer) = 0.40+0.25-0.15= 50%
Therefore
P(no speed or beer) = 1-50% = 50%
==================
Cheers,
Stan H.